A diffraction grating produces a third-order bright spot for wavelength 550 nm at 50o from the central maximum. What is/are the other angle(s) where the bright spot can occur? [angle is expressed in 2 significant figures]

A. 8o

B. 18o

C. 27o

D. 50o

1
Expert's answer
2021-07-19T09:35:44-0400

"\text{For a diffraction grating}\newline
d\sin\theta_m = m\lambda"

"\lambda= 550 nm"

"\theta_3 = 50\degree"

"d= \frac{m\lambda}{\sin\theta_m}=\frac{3*550}{\sin50\degree}\approx2154\ nm"

"m= \frac{d\sin\theta_m}{\lambda}"

"A.\text{if }\theta_m= 8\degree\newline
m= \frac{d\sin8\degree}{\lambda}= \frac{2154\sin8\degree}{550}=0.54\newline
m\notin Z"

"m= \frac{d\sin8\degree}{\lambda}= \frac{2154\sin8\degree}{550}=0.54"

"B.\text{if }\theta_m=18\degree\newline
m= \frac{d\sin18\degree}{\lambda}= \frac{2154\sin18\degree}{550}=1.21\newline
m\notin Z"

"C.\text{if }\theta_m= 27\degree\newline
m= \frac{d\sin27\degree}{\lambda}= \frac{2154\sin27\degree}{550}=1.77\newline
m\notin Z"

"D.\text{if }\theta_m= 50\degree\newline
m= \frac{d\sin50\degree}{\lambda}= \frac{2154\sin50\degree}{550}=3\newline
m\in Z"

"\text{Answer: D}.50\degree"